# Relativistic Axion with Nonrelativistic Momenta: A Robust Bound on Minimal ALP Dark Matter

### v1

@@ Abstract
where $m_{\phi}$ and $f_{\phi}$ denote the ALP mass and decay constant, respectively. This bound is known for the ALP dark matter dominated by the homogeneous zero‐momentum mode, under the requirement that coherent oscillations begin early enough to satisfy the late‐forming dark matter constraint. One loop hole to evade this limit may be to introduce a large amount of the non-relativistic modes of the ALP with non-vanishing momenta. Here we show that the same limit remains valid even if nonzero‐momentum modes dominate. Interstingly, when nonrelativistic gradient modes prevail, the ALP behaves relativistic radiation rather than matter, if it violates the limit. Moreover, if the typical momentum is sufficiently small, domain walls form.
@@ I Introduction
Dark matter provides direct evidence for physics beyond the Standard Model, yet its nature remains unknown. Among the proposed candidates, the axion is particularly well motivated. In the misalignment mechanism Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983), the axion begins coherent oscillations when its mass becomes comparable to the Hubble parameter, and the resulting energy density can account for dark matter. It has also been shown that if inflation lasts sufficiently long and the Hubble parameter during inflation is lower than the QCD scale, the predicted axion abundance can be significantly altered Graham and Scherlis (2018); Takahashi et al. (2018); Ho et al. (2019); Alonso-Álvarez et al. (2020). Additionally, the axion/ALP can be initially placed near the hilltop of its potential through interactions with other fields, thereby delaying the onset of oscillations and enhancing its abundance Daido et al. (2017); Co et al. (2019); Takahashi and Yin (2019); Nakagawa et al. (2020); Narita et al. (2023). Light ALPs can also be produced via stimulated emission, such as from inflaton, Higgs decay or even thermal scattering Moroi and Yin (2021a, b); Nakayama and Yin (2021); Yin (2023); Sakurai and Yin (2025), or through wide resonance effects Co et al. (2018); Harigaya and Leedom (2020).
@@
This potential is bounded above and below, which gives an constraint on ALP dark matter from the onset of coherent oscillations, as shown in, e.g., Refs. Marsh and Yin (2021); Nakagawa et al. (2023) by taking account of constraints for the small scale structure from Lyman-$\alpha$ and Milky Way subhalo count.
@@ II Relativistic axion with non-relativistic momenta
Here we perform a lattice simulation by using Cosmolattice Figueroa et al. (2021, 2023) to study the impact on the non-homogeneous mode given that the potential Eq.(1) is a cosine term, which is non-linear in $\phi$.
@@
$\phi^{0}=\bar{\phi}^{0}+\delta\phi^{0},$ (5)
where $\bar{\phi}^{0}$ is the homogeneous mode, that we take $\bar{\phi}^{0}=0,\dot{\bar{\phi^{0}}}=0$, since the ALP limit on the dark matter with the dominated homogeneous mode is known. $\dot{X}\equiv dX/dt$ with $t$ being the cosmic time. $\delta\phi$ represents the fluctuation. Throughout this paper, we use superscript 00 to denote the initial condition for the simulation.
The fluctuation $\delta\phi^{0}$ satisfies
$\langle\delta\phi^{0}\delta\phi^{0}\rangle=\int^{K_{\rm UV}}_{K_{\rm IR}}\frac{dk}{k}\,{\cal P}_{\delta\phi}(k).$ (6)
Here $K_{\rm UV,IR}$ is the momentum cutoff for the spectrum we consider. We take $K_{\rm IR}\gtrsim H^{0}$, with $H$ being the Hubble parameter, because we consider the fluctuation is adiabatic for conservativeness, and we choose the uniform density slicing for the super-horizon mode. Otherwise we will have constraint from isocurvature modes.
@@
$\langle\delta\phi^{0}\delta\phi^{0}\rangle=\frac{\tilde{\phi}^{2}}{q}\left((\frac{K_{\rm UV}}{m_{\phi}})^{q}-(\frac{K_{\rm IR}}{m_{\phi}})^{q}\right),$ (8)
@@
$\ddot{\phi}-\frac{1}{a^{2}}\Delta\phi_{i}+3H\dot{\phi}=-\partial_{\phi}V_{{\rm eff}},$ (9)
@@
by assuming a radiation dominated Universe for our purpose. $a$ is the scale factor, which is taken to be unity, $a^{0}=1$, at the initial simulation time.
In Fig.1 we show the solution for the energy densities normalized by $a^{3}$. We take $\tilde{\phi}=f_{\phi},K_{\rm UV}=m_{\phi}$ (left panel), $\tilde{\phi}=10f_{\phi},K_{\rm UV}=m_{\phi}$ (middle panel) and $\tilde{\phi}=40f_{\phi},K_{\rm UV}=1/2m_{\phi}$ (right panel). In these panels, we take $H^{0}=0.5m_{0},m_{\phi}^{2}=5m_{0}^{2},q=2,~{}{\rm and}~{}K_{\rm IR}=H_{0}$. Here $m_{0}$ is the machine unit.
Fig. 1: The (conformal) time evolution of the energy density normalized by $a^{3}$. The purple, green, blue and yellow lines denote the kinetic energy, gradient energy, potential energy, and the total energy, respectively. In all panels, we take $H^{0}=0.5m_{0},m_{\phi}^{2}=5m_{0}^{2},q=2,~{}{\rm and}~{}K_{\rm IR}=H_{0}$. In the left, middle and right panels, we choose $(\tilde{\phi},K_{\rm UV})=(f_{\phi},m_{\phi})$, $(10f_{\phi},m_{\phi})$ and $(40f_{\phi},1/2m_{\phi})$ , respecively.
Fig. 2: The histogram of $\phi/f_{\phi}$ at $\tau=8/m_{0}$.
@@
The left panel shows the case that Eq.(3) is satisfied. In the other two panels, Eq.(3) is violated. This can be seen from the initial gradient energy density,
$\left\langle{(\nabla\phi)^{2}}\right\rangle=\tilde{\phi}^{2}\frac{\frac{K_{\rm UV}^{2+q}}{m_{\phi}^{q}}-\frac{K_{\rm IR}^{2+q}}{m_{\phi}^{q}}}{2+q},$ (10)
which is
$\left\langle{(\nabla\phi)^{2}}\right\rangle=\frac{1}{4}\frac{\tilde{\phi}^{2}K_{\rm UV}^{4}}{m_{\phi}^{2}}$ (11)
In the parameter choice for the figure, we have $\frac{1}{4}m_{\phi}^{2}f_{\phi}^{2},25m_{\phi}^{2}f_{\phi}^{2},25m_{\phi}^{2}f_{\phi}^{2}$ from the left to right panels.
@@
We can clearly see that the total energy of the left panel behaves as matter, while the ones in the other two behave more like radiation at the beginning. The initial potential energy in the left panel scales as matter in time while the other two behaves as cosmological constant, i.e. $\left\langle{V}\right\rangle a^{3}\propto a^{3}$, at the beginning and later they behave as matter.
@@
The difference can be understood intuitively as follows. For the relevant mode $k_{\rm IR}<k<k_{\rm UV}\leq m_{\phi}$ with large amplitudes the right-hand side of Eq.(9), which is bounded by $m_{\phi}^{2}f_{\phi}^{2}$ can be neglected. Then the gradient energy becomes comparable to the kinetic energy and ALP behaves as radiation as the massless limit. The potential energy by averaging the field fluctuations in this limit is $\left\langle{V}\right\rangle=m_{\phi}^{2}f_{\phi}^{2}$, which behaves as cosmological constant.
When the gradient/kinetic energy redshifts to the potenial energy, we have two possibilities. When the amplitude is not very large, $m_{\phi}$ is comparable to the typical momentum, one can have the ALP as the matter. This corresponds to the middle panel.
When the typical momentum becomes smaller than the mass, i.e., when the field amplitude is much larger than $f_{\phi}$, the field configurations in coordinate space begin to cross the potential barriers multiple times after the kinetic and gradient energies become comparable to the potential energy. As a result, domain walls are formed. This behavior can be observed in the right panel of Fig. 1, where the gradient energy increases.
Due to a population bias among the vacua, we expect that the domain walls will collapse within a relatively short lifetime, provided the field amplitude is not extremely larger than $f_{\phi}$. The histogram of $\phi$ corresponding to the right panel of Fig. 1 is shown in Fig. 2 at $\tau=8/m_{0}$. It clearly indicates that the field values are localized around multiple vacua $\phi=2n\pi f_{\phi}$ with $n$ being an integer, confirming that domain walls are formed by field configurations connecting these vacua.
@@
One cannot get the matter dominant epoch if kinetic energy/gradient energy is larger than the potential energy at $z<5.5\times 10^{6}$ even if the typical momentum is smaller than the mass. This shows again Eq.(4).
III Conclusions

### v2

@@ Abstract
where $m_{\phi}$ and $f_{\phi}$ denote the ALP mass and decay constant, respectively. This bound is known for the ALP dark matter dominated by the homogeneous zero‐momentum mode, under the requirement that coherent oscillations begin early enough to satisfy the late‐forming dark matter constraint. One loop hole to evade this limit may be to introduce a large amount of the non-relativistic modes of the ALP with non-vanishing momenta. Here we show that the same limit remains valid even if nonzero‐momentum modes dominate. Interestingly, when nonrelativistic gradient and kinetic modes prevail, the ALP behaves relativistic radiation rather than matter, if it violates the limit. Moreover, if the typical momentum is sufficiently small, Baumkuchen-like domain walls form, which play an important role in understanding the transition.
preprint: TU-1268
@@ I Introduction
Dark matter provides direct evidence for physics beyond the Standard Model, yet its nature remains unknown. Among the proposed candidates, the axion is particularly well motivated. In the misalignment mechanism Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983), the axion begins coherent oscillations when its mass becomes comparable to the Hubble parameter, and the resulting energy density can account for dark matter. It has also been shown that if inflation lasts sufficiently long and the Hubble parameter during inflation is lower than the QCD scale, the predicted axion abundance can be significantly altered Graham and Scherlis (2018); Takahashi et al. (2018); Ho et al. (2019); Alonso-Álvarez et al. (2020). Additionally, the axion/ALP can be initially placed near the hilltop of its potential through interactions with other fields, thereby delaying the onset of oscillations and enhancing its abundance Daido et al. (2017); Co et al. (2019); Takahashi and Yin (2019); Nakagawa et al. (2020); Narita et al. (2023). Light ALPs can also be produced via stimulated emission, such as from inflaton, Higgs decay or even thermal scattering Moroi and Yin (2021a, b); Nakayama and Yin (2021); Choi et al. (2023); Yin (2023); Sakurai and Yin (2025), or through wide resonance effects Co et al. (2018); Harigaya and Leedom (2020).
@@
This potential is bounded above and below, which gives an constraint on ALP dark matter from the onset of coherent oscillations, as shown in, e.g., Refs. Marsh and Yin (2021); Dror and Leedom (2020); Nakagawa et al. (2023) by taking account of constraints for the small scale structure from Lyman-$\alpha$ and Milky Way subhalo count.
@@ II Relativistic axion with non-relativistic momenta
Here we perform a lattice simulation by using Cosmolattice Figueroa et al. (2021, 2023) to study the impact on the non-homogeneous mode given that the potential Eq.(1) is a cosine term, which is non-linear in $\phi$. We will show when Eq.(3) is satisfied, even if the momentum is smaller than $m_{\phi}$ the ALP behaves as radiation.
@@
$\phi_{i}=\bar{\phi}_{i}+\delta\phi_{i},$ (5)
where $\bar{\phi}_{i}$ is the homogeneous mode, that we take $\bar{\phi}_{i}=0,\dot{\bar{\phi_{i}}}=0$, since the ALP limit on the dark matter with the dominated homogeneous mode is known. $\dot{X}\equiv dX/dt$ with $t$ being the cosmic time. $\delta\phi$ represents the fluctuation. Throughout this paper, we use the subscript $i$ to denote the initial condition for the simulation.
The fluctuation $\delta\phi_{i}$ satisfies
$\langle\delta\phi_{i}\delta\phi_{i}\rangle=\int^{K_{\rm UV}}_{K_{\rm IR}}\frac{dk}{k}\,{\cal P}_{\delta\phi}(k).$ (6)
Here $K_{\rm UV,IR}$ is the momentum cutoff for the spectrum we consider. Namely, we assume the reduced power spectrum vanish out of the range $K_{\rm IR}-K_{\rm UV}]$. We take $K_{\rm IR}\gtrsim H_{i}$, with $H$ being the Hubble parameter, because we consider the fluctuation is adiabatic for conservativeness, and we choose the uniform density slicing for the super-horizon mode. This is the case for instance the ALP is produced from inflaton decay. In this case, we will not have the constraint from large scale isocurvature modes.
To ensure a vacuum-like state with minimal energy, the field fluctuation and its time derivative are related by the on-shell condition,
$\delta\dot{\phi}_{i,\vec{k}}=\left(\pm i\omega_{k}-aH\right)\,\delta\phi_{i,\vec{k}},$ (7)
where $\delta\phi_{i,\vec{k}}$ and $\delta\dot{\phi}_{i,\vec{k}}$ denote the Fourier modes of $\delta\phi$ and $\delta\dot{\phi}$, respectively. The frequency is given by
$\omega_{k}=\sqrt{\vec{k}^{\,2}+m_{\phi}^{\,2}},$ (8)
following the standard vacuum prescription used in lattice simulations such as CosmoLattice. Here $m_{\phi}$ is the curvature of the potential evaluated at the homogeneous background $\bar{\phi}$; in our setup we take $\bar{\phi}=0$. This initialization corresponds to the usual dispersion relation for ALP “particles”. Simulations with a different choice for the initial condition are shown in Appendix C.
@@
$\langle\delta\phi_{i}\delta\phi_{i}\rangle=\frac{\tilde{\phi}^{2}}{q}\left((\frac{K_{\rm UV}}{m_{\phi}})^{q}-(\frac{K_{\rm IR}}{m_{\phi}})^{q}\right),$ (10)
@@
$\ddot{\phi}-\frac{1}{a^{2}}\Delta\phi+3H\dot{\phi}=-\partial_{\phi}V,$ (11)
@@
by assuming a radiation dominated Universe for our purpose. $a$ is the scale factor, which is taken to be unity, $a_{i}=1$, at the initial simulation time. $\Delta$ is the Laplacian for the comoving coordinate.
Fig. 1 shows the time evolution of the energy density normalized by $a^{3}$. The red dot-dashed, blue dashed, green dotted, and black solid lines represent the kinetic, gradient, potential, and total energy densities, respectively. In all panels we set $H_{i}=0.5\,m_{\star}$, $q=2$, and $K_{\rm IR}=H_{i}$. The five panels correspond to different choices of $(\tilde{\phi},K_{\rm UV})$, namely $(f_{\phi},m_{\phi})$, $(10f_{\phi},m_{\phi})$, $(40f_{\phi},m_{\phi})$, $(40f_{\phi},0.5m_{\phi})$, and $(40f_{\phi},\simeq 0.84m_{\phi})$, from the top-left to the bottom-right. In all panels we take $m_{\phi}^{2}=5m_{\star}^{2}$, except for the bottom-right one, where $m_{\phi}^{2}=10m_{\star}^{2}$. Here $m_{\star}$ denotes the machine unit. [^1] For other parameter sets, our conclusions remain qualitatively unchanged. For 2<q<0-2<q<0, the results differ quantitatively but not qualitatively.
Fig. 1: Time evolution of the energy density normalized by $a^{3}$. The red dot-dashed, blue dashed, green dotted, and black solid lines represent the kinetic, gradient, potential, and total energy densities, respectively. In all panels we set $H_{i}=0.5\,m_{\star}$, $q=2$, $K_{\rm IR}=H_{i}$, and a comoving box size $L\simeq 63/m_{\star}$ with $a_{i}=1$. The top-left, top-middle, top-right, bottom-middle, and bottom-right panels correspond to $(\tilde{\phi},K_{\rm UV})=(f_{\phi},m_{\phi})$, $(10f_{\phi},m_{\phi})$, $(40f_{\phi},m_{\phi})$, $(40f_{\phi},0.5m_{\phi})$, and $(40f_{\phi},\approx 0.84m_{\phi})$, respectively. In all panels we take $m_{\phi}^{2}=5m_{\star}^{2}$, except for the bottom-right one, where $m_{\phi}^{2}=10m_{\star}^{2}$. The bold black segments are added for visual reference to check the scaling behavior. The purple star in the bottom-middle panel indicates the point corresponding to the plots shown in Figs. 2 and 3.
@@
The top-left panel shows the case in which Eq.(3) is satisfied, whereas in the other four panels Eq.(3) is violated. This can be understood from the initial gradient energy density,
$\frac{1}{2}\left\langle{(\nabla\phi_{i})^{2}}\right\rangle=\frac{1}{2}\tilde{\phi}^{2}\frac{K_{\rm UV}^{2+q}/m_{\phi}^{q}-K_{\rm IR}^{2+q}/m_{\phi}^{q}}{2+q},$ (12)
which reduces to
$\frac{1}{2}\left\langle{(\nabla\phi_{i})^{2}}\right\rangle\sim\frac{1}{8}\frac{\tilde{\phi}^{2}K_{\rm UV}^{4}}{m_{\phi}^{2}}.$ (13)
The kinetic energy is estimated
$\frac{1}{2}\left\langle{(\dot{\phi}_{i})^{2}}\right\rangle\sim\frac{m_{\phi}^{2}}{2}\frac{\tilde{\phi}^{2}}{q}\left((\frac{K_{\rm UV}}{m_{\phi}})^{q}-(\frac{K_{\rm IR}}{m_{\phi}})^{q}\right)\sim\frac{1}{4}\tilde{\phi}^{2}K^{2}_{\rm UV},$ (14)
where we used that $K_{\rm UV}\lesssim m_{\phi},$ and approximated $\omega\approx m_{\phi}.$ This is comparable to or larger than the gradient energy when $K_{\rm UV}\lesssim m_{\phi}.$
For the parameter choices in Fig. 1, the initial values of $\left\langle{(\nabla\phi_{i})^{2}}\right\rangle$ are $\frac{1}{4}m_{\phi}^{2}f_{\phi}^{2}$, $25m_{\phi}^{2}f_{\phi}^{2}$, $400m_{\phi}^{2}f_{\phi}^{2}$, $25m_{\phi}^{2}f_{\phi}^{2}$, and $400m_{\phi}^{2}f_{\phi}^{2}$, from the top-left to the bottom-right panels, respectively. These values are also indicated at the top of each panel. [^2] The simulations in the top panels employ $512^{3}$ lattices, while those in the bottom panels use $1024^{3}$ lattices.
@@
We clearly see that the total energy in the top-left panel behaves as matter, whereas in the other panels it initially behaves like almost radiation. The potential energy in the top-left panel scales as matter, while in the other cases it first behaves as a cosmological constant, $\left\langle{V}\right\rangle a^{3}\propto a^{3}$, and later transitions to matter-like behavior.
@@
This difference can be intuitively understood as follows. For the relevant modes $k_{\rm IR}<k<k_{\rm UV}\leq m_{\phi}$ with large amplitudes, the right-hand side of Eq.(11), bounded by $m_{\phi}^{2}f_{\phi}$, can be neglected. In this regime, the gradient energy becomes comparable to the kinetic energy, and the ALP field behaves as radiation, i.e., the massless limit. The averaged potential energy is then $\left\langle{V}\right\rangle=m_{\phi}^{2}f_{\phi}^{2}$, which effectively behaves as a cosmological constant.
As the gradient and kinetic energies redshift and become comparable to the potential energy, a transition in the equation of state occurs in the top-middle to bottom-right panels. Analytically, this transition happens at $a\sim 2.5$, $4.9$, $3.3$, and $4.4$, for the top-middle, top-right, bottom-middle, and bottom-right, respectively. [^3] For discussing the limit in Eq.(2), this period can be regarded as corresponding to $z_{\rm matter}$. At this stage, $m_{\phi}$ is about an order of magnitude larger than $H$. Although increasing this hierarchy is computationally challenging in lattice simulations, we expect that our results would not change significantly for a larger hierarchy, as can be inferred from the comparison between the top-right and bottom-right panels in Fig. 1. The figure indeed shows that the behavior changes at these epochs, particularly in the potential energy. By comparing the top and bottom panels, one can see that the transition epoch depends only weakly on $K_{\rm UV}$, $\tilde{\phi}$, or $m_{\phi}$, as long as the initial total energy density is fixed.
Fig. 2: Distribution of $\phi/f_{\phi}$ for the bottom-middle panel of Fig. 1. The lower panel focuses on the short time interval until $a=3$ (c.f. the transition at $a\sim 2.5$).
Fig. 3: Snapshots of $\phi/f_{\phi}$ on an $x$-$y$ slice (we choose the 513th slice along the $z$ direction) at $a=1,3,5,7,9,$ and $11$. The black contours in the top panels mark the locations of the domain walls.
Relation of transition to formation of Baumkuchen domain wall
Note that when this transition occurs with nonrelativistic momenta, $H<k_{\rm typ}<m_{\phi}$, the typical field amplitude $|\delta\phi_{\rm amp}|$ satisfying $\left\langle{(\nabla\phi)^{2}}\right\rangle\sim k_{\rm typ}^{2}\delta\phi_{\rm amp}^{2}\sim m_{\phi}^{2}f_{\phi}^{2}$ is larger than $f_{\phi}$. Thus, the field configurations in coordinate space cross multiple potential barriers, and domain walls are formed around or slightly before the transition epoch. Due to the population bias among different vacua, we expect the domain walls to collapse within a relatively short lifetime.
The histogram of $\phi$ corresponding to the bottom-middle panel of Fig. 1 is shown in the top panel of Fig. 2 (see also the other simulations in Appendices A and C). It clearly indicates that the field values become localized around multiple vacua, $\phi=2n\pi f_{\phi}$ with integer $n$, soon after $a\gtrsim 2.5$, confirming that domain walls are indeed formed by field configurations connecting these vacua. Subsequently, the domain wall network collapses into a single domain around $\phi/f_{\phi}\simeq 0$, as seen for $a\to 9$.
Together with the bottom panel of Fig. 2, which focuses on the short period $a\lesssim 3$, we can trace the detailed dynamics of this transition. One can observe that the field distribution already exhibits a mild local preference before the transition, implying that the wall configurations start forming even earlier. However, during this stage, the wall energy is negligible compared with the total energy, which is dominated by gradient and kinetic terms.
Indeed, the emerging domain walls exhibit an approximately layered, chain-like structure along the direction perpendicular to the walls, if one neglects their curvature locally. The wall locations are characterized by
$\phi_{\rm DW}/f_{\phi}={\{\cdots,-3\pi,-\pi,\pi,3\pi,\cdots\}}\subset[-\delta\phi_{\rm amp}/f_{\phi},\,\delta\phi_{\rm amp}/f_{\phi}].$ (15)
Note that domain walls can only annihilate with those possessing the same $\phi_{\rm DW}/f_{\phi}$. For illustration, consider an axion wave with wavelength $2\pi/k_{\rm typ}$ in the $z$-direction and amplitude $\delta\phi_{\rm amp}$,
$\phi(x)\propto\delta\phi_{\rm amp}\cos(k_{\rm typ}z),$ (16)
where we neglect the time-dependence. In this case, the field repeatedly crosses adjacent vacua as a function of $z$, and a chain of nearly parallel domain walls is formed. The number of walls within a half wavelength $\pi/k_{\rm typ}$ is approximately $\sim 2\delta\phi_{\rm amp}/(2\pi f_{\phi})$. As long as the gradient and kinetic energies dominate over the potential energy, the wall tension is subdominant compared to the wave energy density, so the walls do not enter a scaling regime but instead approximately follow the underlying axion-wave pattern.
When such a chain of walls closes into a loop, it forms a layered, Baumkuchen-like domain-wall configuration. We refer to these layered loops as “Baumkuchen domain walls”. The energy density of a Baumkuchen domain-wall loop can be estimated by assuming a single wall within a box of size $(\pi/k_{\rm typ})\times 2\pi f_{\phi}/(2\delta\phi_{\rm amp})$. We then obtain
$\rho_{\rm BDW}\sim\frac{8f_{\phi}^{2}m_{\phi}}{2\pi^{2}f_{\phi}/(2k_{\rm typ}\delta\phi_{\rm amp})}=\frac{8}{\pi^{2}}f_{\phi}m_{\phi}\,\delta\phi_{\rm amp}\,k_{\rm typ}.$ (17)
This energy density remains smaller than the gradient and kinetic energies as long as $m_{\phi}^{2}f_{\phi}^{2}\lesssim k_{\rm typ}^{2}\delta\phi_{\rm amp}^{2}$.
At the transition ($a\sim 2.5$ in Fig. 2), the energy stored in the Baumkuchen domain walls becomes significant, and the wall tension drives the field distribution toward $\phi=0$, the center of the initial condition. The Baumkuchen domain-wall network then approaches a scaling regime, containing $\mathcal{O}(1)$ large Baumkuchen domain wall per horizon, while smaller domain wall loops collapse. This evolution can be understood as a layered collapse of walls with different $\phi_{\rm DW}$ values. Such behavior can be seen in the snapshots of Fig. 3 for $a=3$ and $a=5$, where we also show the wall positions by the black contours. The longest surviving walls are associated with the domains containing the most populated vacuum, while smaller internal walls collapse earlier. The largest walls cannot collapse before the smaller ones due to the interaction forces between walls with different $\phi_{\rm DW}$.
Then the domain walls with no superhorizon correlation (see Appendix B) collapse shortly thereafter, completing the transition. After this collapse, axion particles are produced from the decay of the domain walls and subsequently behave as ordinary cold matter. At the moment of production, these particles are semi-relativistic, with typical momenta of order $m_{\phi}$, originating from the dynamics of the domain-wall collapse. This completes the overall picture of particle production and energy transfer associated with this nonlinear transition.
We also observe in the bottom panel of Fig. 2 that, before the formation of the domain walls, the field distribution becomes broader and exhibits a characteristic temporal oscillation. This originates from our “nonrelativistic particle–like” initial condition, in which the initial kinetic energy is slightly larger than the gradient energy. When the initial kinetic energy dominates, the positive- and negative-frequency components of each mode acquire almost the same initial amplitude, and their interference produces a standing-wave component. This leads to an oscillatory exchange between the kinetic and gradient energies, as clearly visible in the four right panels of Fig. 1. The amplitude of this oscillation is gradually reduced by cosmological redshift. In this sense, the initially large kinetic energy drives the system into a regime where the kinetic and gradient contributions become balanced on time average. When we impose an alternative, physically motivated initial condition that corresponds to a massless “propagating wave” (i.e. $\omega=k$), the left- and right-moving components no longer have correlated phases. As a result, the standing-wave contribution is strongly suppressed and the oscillatory interference pattern disappears. A detailed discussion and analysis as well as numerical simulation is provided in Appendix C, and the phenomena found in the main part hold as well. What we have shown is that, even if we impose an initial condition corresponding to nonrelativistic axion particles, the subsequent evolution behaves like a standing-wave radiation fluid and inevitably leads to the formation of Baumkuchen domain walls.
@@
One cannot obtain a matter-dominated epoch if the kinetic and gradient energies remain larger than the potential energy at redshifts $z<5.5\times 10^{6}$, even when the typical momentum is smaller than the mass. This once again confirms the condition given in Eq.(4).
III Conclusions and discussion